OpenStudy (anonymous):

integrate ∫[dx/(2x)(3x^2+4),x]

OpenStudy (anonymous):

\[\int \frac{1}{2x(3x^2 + 4)}\]?

OpenStudy (anonymous):

∫dx/(2x)(3x^2+4)

OpenStudy (anonymous):

\[\int\limits \left(\begin{matrix}dx \\(2x)(3x^2+4)\end{matrix}\right)\]